# Beyond PCA: Manifold Dimension Estimation via Local Graph Structure

### v1

@@ Abstract
Local principal component analysis (Local PCA) has proven to be an effective tool for estimating the intrinsic dimension of a manifold. More recently, curvature-adjusted PCA (CA-PCA) has improved upon this approach by explicitly accounting for the curvature of the underlying manifold, rather than assuming local flatness. Building on these insights, we propose a general framework for manifold dimension estimation that captures the manifold’s local graph structure by integrating PCA with regression-based techniques. Within this framework, we introduce two representative estimators: quadratic embedding (QE) and total least squares (TLS). Experiments on both synthetic and real-world datasets demonstrate that these methods perform competitively with, and often outperform, state-of-the-art alternatives.
@@ 1 Introduction
Traditional dimensionality reduction techniques, such as principal component analysis (PCA, Wold et al. (1987)), typically assume that data points lie in a subspace of the ambient space $\mathbb{R}^{p}$. Hence, there exists a global linear map that provides a low-dimensional representation of the entire dataset.
@@
A natural research problem arising from the manifold hypothesis is to estimate the intrinsic dimension $d$ from data—a task commonly referred to as manifold dimension estimation. Numerous methods have been proposed across a wide range of domains (e.g., Verveer and Duin (1995); Kégl (2002), Yang et al. (2007); Carter et al. (2009); Sricharan et al. (2010), Kalantana and Einbecka (2013); Facco et al. (2017); Lim et al. (2024)), see Bi and Lafaye de Micheaux (2025) for the most recent survey on this topic. Broadly speaking, most manifold dimension estimators are built upon the so-called flatness assumption, which approximates local regions of the underlying manifold as $d$-dimensional linear spaces. This assumption transforms the inherently nonlinear problem into a linear one, enabling the design of numerous estimators.
@@
In the survey, we conducted an extensive set of experiments to benchmark the performance of eight representative manifold dimension estimators. Among them, three emerged as top performers: two are flatness-based—namely, the dimensionality estimator based on angle and norm concentration (DanCo, Ceruti et al. (2014)) and the two-nearest-neighbor estimator (TwoNN, Facco et al. (2017)); the third is the curvature-adjusted PCA estimator (CA-PCA, Gilbert and O’Neill (2025)).
@@
CA-PCA explicitly incorporates curvature information into the local PCA procedure, thereby improving upon the already competitive, though often overlooked, Local PCA estimator (Fukunaga and Olsen, 1971). TwoNN pushes the flatness assumption to its extreme, mitigating curvature effects by relying solely on the two nearest neighbors of each point. DanCo is based on the distribution of distances and angles among data points within neighborhoods, assuming local flatness, and further improves robustness in high-dimensional settings through a Kullback–Leibler (KL) divergence–based comparison framework.
@@
Despite their superior performance relative to other estimators, as discovered in the survey, there remains substantial room for improvement: DanCo is particularly time-consuming; TwoNN tends to underestimate when the intrinsic dimensionality is high; and CA-PCA may overestimate the dimension when the manifold is nonlinearly embedded in an ambient space of much higher dimension.
@@
Motivated by the strengths and limitations of existing estimators, particularly the robustness of PCA-based methods and the improvements of CA-PCA gained by explicitly accounting for curvature—we propose a framework for designing manifold dimension estimators that recovers the local graph structure of the manifold via regression on PCA coordinates extracted from each neighborhood. This approach enables flexible modeling of nontrivial manifold geometry and noise by embedding the problem into a regression setting, rather than attempting to compute geometric quantities directly from sparse data. As a result, it supports more stable and accurate dimension estimation.
Two representative estimators within this framework are described in details: the quadratic embedding estimator (QE) and the total least squares estimator (TLS). Our experiments show that these estimators—despite their simplicity—achieve performance comparable to state-of-the-art methods—and in many cases, especially with small sample sizes or nonlinear embeddings, outperform them.
@@ 2 Manifold Geometry
In this section, we briefly review several fundamental manifold constructions essential for understanding the methods introduced in this article, before presenting our general design framework for manifold dimension estimation. A more comprehensive treatment of the theoretical foundations is provided in our previous survey. For readers interested in more advanced topics in differential geometry, we refer to Lee (2003) and Lee (2018), which offer detailed expositions of the relevant mathematical background.
@@ 2.1 Tangent Space and Local Graph Representation
The tangent space approximation forms the basis of the widely used flatness assumption in manifold dimension estimation. Suppose we are given a point $\bm{x}_{k}$ along with its $K$-nearest neighbors sampled from $M$. The flatness assumption ignores any geometric or distributional irregularities in the neighborhood and posits the data are uniformly distributed within a $d$-dimensional ball $B_{\bm{x}_{k}}(R)\subseteq T_{\bm{x}_{k}}M$. This local linearization enables the construction of various statistics related to $d$. In this context, “flatness” refers to ignoring curvature at a local scale.
(a) Local representation around $\bm{x}_{0}$.
@@
Furthermore, the neighborhood around $\bm{x}_{0}$ can be reparameterized to better reflect the manifold’s local geometry. By translating $\bm{x}_{0}$ to the origin, we construct a new local coordinate system consisting of $d$ orthonormal vectors spanning the tangent space $T_{\bm{x}_{0}}M$, and pdp-d orthonormal vectors spanning the normal space $T_{\bm{x}_{0}}M^{\perp}$, as shown in Figure 2(a). Standard results from differential geometry (Lee, 2003) then guarantee that any point $\bm{x}$ sufficiently close to $\bm{x}_{0}$ can be expressed in these coordinates as
@@
where $\bm{x}^{\prime}_{a:b}$ denotes the $a$th to $b$th components of $\bm{x}^{\prime}$ (inclusive), $\bm{x}^{\prime}_{0}=0$, and $g\colon\mathbb{R}^{d}\to\mathbb{R}^{p-d}$ is a smooth function satisfying $\nabla g(0)=0$.
@@
In this new coordinate system, the neighborhood of $\bm{x}_{0}$ is represented as the graph of a function $g$. The first $d$ components, $\bm{x}^{\prime}_{1:d}$, serve as coordinates in the tangent space, while the remaining pdp-d components are determined by $g$, capturing deviations from the tangent space approximation. The flatness assumption corresponds to the case where $g\doteq 0$, meaning the last pdp-d coordinates vanish—equivalent to retaining only the first-order Taylor expansion of $g$ around the origin. Consequently, the nontrivial local geometry of the manifold are entirely encoded in the higher-order terms of the Taylor expansion of $g$. In particular, the second-order Taylor polynomial of $g$ provides a quadratic approximation, as illustrated in Figure 2(b).
@@ 2.3 Capturing Local Graph Structure
Let $\mathcal{M}$ denote the set of all manifolds embedded in $\mathbb{R}^{p}$. A natural way to partition $\mathcal{M}$ is by classifying its elements according to their intrinsic dimensions. Alternatively, since any $d$-dimensional manifold can locally be represented as the graph of a smooth function with $d$-dimensional inputs and (pd)(p-d)-dimensional outputs — which we refer to as a $C_{d}^{\infty}$ function — the components of this partition can equivalently be expressed as
$\mathcal{M}_{j}=\left\{\,M\in\mathcal{M}:M\text{ is locally the graph of }C_{j}^{\infty}\text{ functions}\,\right\},$
where $1\leq j\leq p$.
Given a dataset in $\mathbb{R}^{p}$ that satisfies the manifold hypothesis, the underlying manifold must belong to one of $\mathcal{M}_{j}$s. Our approach to manifold dimension estimation is based on recovering the correct local graph structure of this manifold across neighborhoods — that is, identifying the appropriate function class $C_{j}^{\infty}$ from which the local representation functions belong, which can be approached by fitting the data to candidate functions specified for different $\mathcal{M}_{j}$s.
More specifically, let $\{G_{j}\}_{j=1}^{p}$ denote our full candidate model specification, where each $G_{j}$ is a family of functions from $C_{j}^{\infty}$ we choose to model local graph structures of a $j$-dimensional manifold, i.e., $G_{j}$ corresponds to certain classes of $j$-dimensional manifolds. For each candidate dimension $j$, the data is fitted to the corresponding model class $G_{j}$, and a goodness-of-fit score $\sigma_{j}$ is computed to assess the quality of the fit. Ideally, the model is (approximately or exactly) correctly specified only at the true intrinsic dimension $j=d$, which should yield the best fit and, consequently, the highest score. Therefore, a dimension estimator can be constructed as
@@
Further details about the model fitting procedure and evaluation procedure will be discussed in the next section.
In its most general form, a full model specification may span all $p$ possible partition classes. Within each partition, one may consider multiple families of $C_{j}^{\infty}$ functions, each tailored to model specific local regions of the manifold. The theory guarantees the existence of local graph representation under very mild conditions: it suffices for $g\in C_{j}^{\infty}$ to be a smooth function satisfying $\nabla g(0)=0$. However, this level of generality is typically too broad to support practical estimation or meaningful model comparison.
To make progress, we must constrain the function classes used to fit data within each partition. The choice of these classes should ideally be guided by our prior knowledge of the underlying manifold. In the absence of such information, a natural default is to take each $G_{j}$ as a family of quadratic functions. Specifically, we approximate the output of the graph function $g\colon\mathbb{R}^{j}\to\mathbb{R}^{p-j}$ using a tensor of quadratic forms. Let $\mathcal{Q}\in\mathbb{R}^{(p-j)\times j\times j}$ be a third-order tensor, where each slice $\mathcal{Q}[\ell,:,:]=Q_{\ell}$ is a symmetric matrix encoding the second-order variation of the manifold in the $\ell$-th output direction. Then the functions in $G_{j}$ take the form
@@ 3.1 Regression with PCA Coordinates
Intuitively, and supported by the strong empirical performance of Local PCA—which, at its core, uses the span of principal directions to approximate the tangent space—these two approximations should be close at sufficiently small scales. In other words, the subspace spanned by the first $d$ eigenvectors provides an estimate of the tangent space, up to a small translation and rotation. Indeed, PCA is widely regarded as the most natural method for tangent space estimation; see Lim et al. (2024) for a theoretical discussion of its approximation properties.
For our purposes, the most important consequence is that PCA coordinates serve as good estimates of the theoretical (new) coordinates appearing in the local graph representation, effectively enabling us to fit the models specified for different dimensions. At its core, our estimator seeks to identify the splitting index $d$ at which a deterministic relationship emerges between the first $d$ coordinates and the remaining pdp-d, governed by the local graph function. The correct intrinsic dimension should correspond to a clear improvement in model fit, thereby revealing the true value of $d$.
@@
Our QE estimator adopts this approximation but uses only the first output coordinate (i.e., $x_{k,j+1}$, corresponding to $q_{1}$) as the response. This design has two key advantages: first, PCA coordinates associated with smaller eigenvalues are typically more sensitive to noise; second, it helps avoid spurious relationships between input and output when $j<d$. In contrast, using all of the last pjp-j coordinates as responses introduces the risk of detecting artificial dependencies, which can negatively affect model evaluation.
Our TLS-based estimator (TLS) mirrors the structure of QE, with the key distinction that the regression step is performed using total least squares (TLS) instead of OLS. We also experimented with LASSO (Tibshirani, 1996) and ridge regression (McDonald, 2009) as alternative fitting methods, but found that they yielded inferior performance in this context.
@@ 3.2 Filling Details: QE and TLS
Evaluating the model fit is a crucial step within our framework, as (2) suggests, a good fit signals that the intrinsic dimension has likely been reached. For QE, this step is straightforward: classical statistical theory provides the $F$-statistic, specifically designed to compare a target regression model against a trivial one.
Figure 4 shows the distribution of $F$-statistics for $j=1,2,3$, computed from a sample of size $1000$ drawn uniformly from a $d=3$ deformed sphere (with curvature parameter $c=0.01$). A clear jump in the $F$-statistic is observed at $j=3$, accurately identifying the intrinsic dimension.
@@
Despite this, the $F$-statistic is well-known for its robustness to such violations (Box and Watson, 1962; Ali and Sharma, 1996). Empirically, we consistently observe a pronounced jump in its value at the true dimension, and applying a fixed $p$-value threshold (e.g., at the 1% level) leads to reliable results across a broad range of datasets and settings.
@@
    Since p1p-1 models are fitted in each neighborhood, the procedure can become inefficient when $p\gg d$. In addition to implementation-level optimizations such as parallelization, we limit the number of models per neighborhood to a maximum of $p_{\max}-1$. This threshold is determined by three factors: the ambient dimension $p$; the largest index for which the standard deviation of the new coordinates is sufficiently large (to avoid fitting near-constant responses); and the neighborhood size $K$, as regression becomes unreliable when $K$ is smaller than the number of predictors. Letting $q$ denote the largest nonconstant coordinate index, we define:
    $p_{\max}=\min\left\{p,q,\left\lfloor\sqrt{1+8K}-1\right\rfloor\right\}.$ (5)

### v2

@@ Abstract
Local principal component analysis (Local PCA) has proven to be an effective tool for estimating the intrinsic dimension of a manifold. More recently, curvature-adjusted PCA (CA-PCA) has improved upon this approach by explicitly accounting for the curvature of the underlying manifold, rather than assuming local flatness. Building on these insights, we propose a general framework for manifold dimension estimation that characterizes the manifold’s local graph structure through the integration of PCA and regression-based techniques. Within this framework, we introduce two representative estimators: quadratic embedding (QE) and total least squares (TLS). Experiments on both synthetic and real-world datasets demonstrate that these methods perform competitively with, and often outperform, state-of-the-art alternatives.
@@ 1 Introduction
Traditional dimensionality reduction techniques, such as principal component analysis (PCA, Wold et al. (1987)), typically assume that data points lie in a subspace of the ambient space $\mathbb{R}^{p}$. Under this assumption, there exists a global linear map that provides a low-dimensional representation of the entire dataset.
@@
A natural research problem arising from the manifold hypothesis is to estimate this intrinsic dimension $d$ from data—a task commonly referred to as manifold dimension estimation. Numerous methods have been proposed across a wide range of domains over the years; see, e.g., Verveer and Duin (1995), Kégl (2002), Yang et al. (2007), Carter et al. (2009), Sricharan et al. (2010), Kalantana and Einbecka (2013), Facco et al. (2017), Lim et al. (2024), and Bi and Lafaye de Micheaux (2025) for a recent survey and numerical comparisons on this topic. Broadly speaking, most manifold dimension estimators are built upon the so-called flatness assumption, which approximates local regions of the underlying manifold as $d$-dimensional linear spaces. This assumption transforms the inherently nonlinear problem into a linear one, enabling the design of numerous estimators.
@@
In Bi and Lafaye de Micheaux (2025), an extensive set of experiments was conducted to benchmark the performance of eight representative methods for manifold dimension estimation. Among them, three emerged as top performers: two are flatness-based—namely, the dimensionality estimator based on angle and norm concentration (DanCo, Ceruti et al. (2014)) and the two-nearest-neighbor estimator (TwoNN, Facco et al. (2017)); the third is the curvature-adjusted PCA estimator (CA-PCA, Gilbert and O’Neill (2025)).
@@
DanCo estimates the intrinsic dimension by analyzing the distributions of pairwise distances and interpoint angles within neighborhoods under the flatness assumption. It further enhances robustness in high-dimensional settings through a Kullback–Leibler (KL) divergence–based comparison framework. TwoNN pushes the flatness assumption to its extreme, mitigating curvature effects by relying solely on the two nearest neighbors of each point. Both estimators therefore ignore curvature in local regions. In contrast, CA-PCA explicitly incorporates curvature information into the local PCA procedure, thus improving upon the already competitive yet often overlooked Local PCA estimator (Fukunaga and Olsen, 1971).
@@
Despite their superior performance relative to other estimators, as shown by extensive simulations in Bi and Lafaye de Micheaux (2025), there remains substantial room for improvement: DanCo is particularly time-consuming; TwoNN tends to underestimate when the intrinsic dimensionality is high, and CA-PCA may overestimate the dimension when the manifold is nonlinearly embedded in an ambient space of much higher dimension.
@@
Motivated by the strengths and limitations of existing estimators, particularly the robustness of PCA-based approaches and the improvement achieved by CA-PCA through its explicit treatment of curvature—we propose a new regression-based framework for manifold dimension estimation. Instead of directly computing geometric quantities from sparse data, our approach recovers the local graph structure of the manifold by regressing normal components on tangent components extracted through PCA from each neighborhood, enabling flexible modeling of curvature and noise.
Our main contributions are as follows:
  • 1.
    Regression-based formulation. We introduce a new framework that embeds manifold dimension estimation into a regression setting, allowing complex geometry and noise to be modeled without relying on direct geometric computation. To the best of our knowledge, this perspective is novel in the manifold dimension estimation literature.
  • 2.
    Representative estimators. We instantiate the framework with two simple yet effective methods: quadratic embedding (QE), a second-order regression model capturing curvature effects, and total least squares (TLS), which addresses errors in both predictors and responses.
  • 3.
    Improved stability and empirical performance. By reframing dimension estimation as regression, the proposed approach achieves improved robustness and accuracy, particularly in small-sample settings and for manifolds with complex curvature. Extensive experiments on both synthetic and real-world datasets demonstrate that QE and TLS achieve competitive or superior performance compared to state-of-the-art methods.
@@ 2 Manifold Geometry
In this section, we briefly review several fundamental manifold constructions essential for understanding the methods introduced in this article, before presenting our general design framework for manifold dimension estimation. A more comprehensive treatment of the theoretical foundation is provided in Bi and Lafaye de Micheaux (2025). For readers interested in more advanced topics in differential geometry, we refer to Lee (2003, 2018), which offer detailed expositions of the relevant mathematical background.
@@ 2.1 Tangent Space and Local Graph Representation
The tangent space approximation forms the basis of the widely used flatness assumption in manifold dimension estimation. Suppose we are given a data point $\bm{x}_{k}$ along with its $K$-nearest neighbors sampled from $M$. The flatness assumption ignores any geometric or distributional irregularities in the neighborhood and posits the data are uniformly distributed within a $d$-dimensional ball $B_{\bm{x}_{k}}(R)\subseteq T_{\bm{x}_{k}}M$. This local linearization enables the construction of various statistics related to $d$. In this context, “flatness” refers to ignoring curvature at a local scale.
(a) Local coordinates around $\bm{x}_{0}$.
@@
Furthermore, the neighborhood around $\bm{x}_{0}$ can be reparameterized to better reflect the manifold’s local geometry. By translating $\bm{x}_{0}$ to the origin, we can construct a new coordinate system consisting of $d$ orthonormal vectors spanning the tangent space $T_{\bm{x}_{0}}M$, and pdp-d orthonormal vectors spanning the normal space $(T_{\bm{x}_{0}}M)^{\perp}$, as shown in Figure 2(a). Standard results from differential geometry (Lee, 2003) then guarantee that any point $\bm{x}$ sufficiently close to $\bm{x}_{0}$ can be expressed in these new coordinates as
@@
where $\bm{x}^{\prime}_{1:d}=(x^{\prime}_{1},\ldots,x^{\prime}_{d})^{\top}$, $\bm{x}^{\prime}_{0}=\bm{0}$, and $g\colon\mathbb{R}^{d}\to\mathbb{R}^{p-d}$ is a smooth function satisfying $\nabla g(\bm{0})=\bm{0}$.
In this new coordinate system, the neighborhood of $\bm{x}_{0}$ is represented as the graph of $g$. The first $d$ components, $\bm{x}^{\prime}_{1:d}$, serve as coordinates in the tangent space, while the remaining pdp-d components are determined by $g$ and capture deviations from the tangent space approximation. The flatness assumption corresponds to the case where $g\doteq 0$, meaning the last pdp-d coordinates vanish—equivalent to retaining only the first-order Taylor expansion of $g$ around the origin. Consequently, the nontrivial local geometry of the manifold is entirely encoded in the higher-order terms of the Taylor expansion of $g$. In particular, the second-order Taylor polynomial of $g$ provides a quadratic approximation, as illustrated in Figure 2(b).
@@ 2.3 Capturing Local Graph Structure
Let $\mathcal{M}$ denote the set of all manifolds embedded in $\mathbb{R}^{p}$. A natural way to partition $\mathcal{M}$ is to group its elements according to their intrinsic dimensions. Equivalently, since any $j$-dimensional manifold can locally be expressed as the graph of smooth functions with $j$-dimensional inputs and (pj)(p-j)-dimensional outputs—whose class is denoted by $C_{j}^{\infty}$—the set $\mathcal{M}$ can be written as the disjoint union
$\mathcal{M}=\bigsqcup_{j=1}^{p}\mathcal{M}_{j}=\bigsqcup_{j=1}^{p}\left\{M\in\mathcal{M}:M\text{ is locally the graph of }C_{j}^{\infty}\text{ functions}\right\}.$
Given a dataset in $\mathbb{R}^{p}$ that satisfies the manifold hypothesis, the underlying manifold must belong to one of these subsets $\mathcal{M}_{j}$—specifically, to $\mathcal{M}_{d}$ if its intrinsic dimension is $d$. Our approach to manifold dimension estimation aims to recover the correct local graph structure of the underlying manifold across its neighborhoods. In other words, we seek to identify the function class $C_{j}^{\infty}$ from which the local representation functions originate, which best captures the relationships among the components of observed data points.
To operationalize this idea, we specify for each candidate dimension $j$ a model family $G_{j}\subseteq C_{j}^{\infty}$. The data are then fitted to the models in $G_{j}$, and a corresponding goodness-of-fit score $\sigma_{j}$ is computed to quantify how well each model family explains the data. Ideally, the model is correctly specified (or best approximated) when $j=d$, resulting in the highest score. Hence, the intrinsic dimension is estimated as
@@
Details of the model-fitting and evaluation procedures are presented in the next section.
In its most general form, a full model specification may span all $p$ possible partition classes. Within each partition, one may consider multiple families of $C_{j}^{\infty}$ functions, each tailored to model specific local regions of the manifold. The theory guarantees the existence of local graph representation under very mild conditions: it suffices for $g\in C_{j}^{\infty}$ to be a smooth function satisfying $\nabla g(\bm{0})=\bm{0}$. However, this level of generality is typically too broad to support practical estimation or meaningful model comparison.
To make progress, we must constrain the function classes used to fit data within each partition. The choice of these classes should ideally be guided by our prior knowledge of the underlying manifold. In the absence of such information, a natural default is to take each $G_{j}$ as the family of quadratic functions. Specifically, we approximate the output of the graph function $g\colon\mathbb{R}^{j}\to\mathbb{R}^{p-j}$ using a tensor of quadratic forms. Let $\mathcal{Q}\in\mathbb{R}^{(p-j)\times j\times j}$ be a third-order tensor, where each slice $\mathcal{Q}[\ell,:,:]=Q_{\ell}$ is a symmetric matrix encoding the second-order variation of the manifold in the $\ell$-th output direction. Then the functions in $G_{j}$ take the form
@@ 3.1 Regression with PCA Coordinates
Intuitively, and supported by the strong empirical performance of Local PCA—which, at its core, uses the span of principal directions to approximate the tangent space—these two approximations should be close at sufficiently small scales. In other words, the subspace spanned by the first $d$ eigenvectors provides an estimate of the tangent space, up to a small translation and rotation. Indeed, PCA is widely regarded as the most natural method for tangent space estimation; see Lim et al. (2024) for a theoretical discussion of its approximation properties. In subsection 3.3, we also provide a statistical justification of this approximation in the population sense.
For our purposes, the most important consequence is that PCA coordinates provide accurate empirical estimates of the theoretical coordinates involved in the local graph representation (1), thereby allowing us to fit the models corresponding to different candidate dimensions. At its core, our estimator operates by identifying the splitting index $d$ at which a deterministic relationship emerges between the first $d$ coordinates and the remaining pdp-d, governed by the local graph function. The correct intrinsic dimension corresponds to the point where a marked improvement in model fit occurs, thereby revealing the true value of $d$.
@@
Our QE estimator adopts this approximation but uses only the first output coordinate (i.e., $x_{k,j+1}^{\prime}$, corresponding to $q_{1}$) as the response. This design has several advantages: first, PCA coordinates associated with smaller eigenvalues are typically more sensitive to noise; second, it helps avoid spurious relationships between input and output when $j<d$. In contrast, using all of the last pjp-j coordinates as responses introduces the risk of detecting artificial dependencies, which can negatively affect model evaluation; lastly, using all output coordinates can easily lead to a combinatorial explosion in the number of parameters, especially when $p$ is large, rendering the models infeasible to fit with the limited data available in neighborhoods.
Our TLS-based estimator (TLS) mirrors the structure of QE, with the key distinction that the regression step is performed using total least squares (TLS, Golub and Van Loan (1980)) instead of OLS. We also experimented with LASSO (Tibshirani, 1996) and ridge regression (McDonald, 2009) as alternative fitting methods, but found that they yielded inferior performance in this context.
@@ 3.2 Filling Details: QE and TLS
Evaluating the model fit is a crucial step within our framework. As (2) suggests, a good fit signals that the intrinsic dimension has likely been reached. For QE, this step is straightforward: classical statistical theory provides the $F$-statistic, specifically designed to compare a target regression model against others.
Figure 4 shows the distribution of $F$-statistics for $j=1,2,3$, computed from a sample of size $1000$ drawn uniformly from a $d=3$ deformed sphere (with parameter $c=0.01$). A clear shift of the $F$-statistic histogram is observed at $j=3$, demonstrating its potential to accurately identify the intrinsic dimension.
@@
Despite this, the $F$-statistic is well-known for its robustness to different kinds of violations (Box and Watson, 1962; Ali and Sharma, 1996). Empirically, we consistently observe a pronounced jump in its value at the true dimension, and applying a fixed $p$-value threshold (e.g., at the 1% level) leads to reliable results across a broad range of datasets and settings.
@@
    Since p1p-1 models are fitted in each neighborhood, the procedure can become inefficient when $p\gg d$. In addition to implementation-level optimizations such as parallelization across neighborhoods, we limit the number of models per neighborhood to a maximum of $p_{\max}-1$. The threshold $p_{\max}$ is determined by three factors: the ambient dimension $p$; the largest index for which the standard deviation of the new coordinates is sufficiently large (to avoid fitting near-constant responses); and the neighborhood size $K$, as regression becomes unreliable when $K$ is smaller than the number of predictors. Letting $q$ denote the largest nonconstant coordinate index, we define [^1] When $j=p_{\max}-1$ components are used as inputs, we need $(p_{\max}-1)(p_{\max}-2)/2+2(p_{\max}-1)+1\leq K+1$, which leads to the expression in (5).:
    $p_{\max}=\min\left\{p,q,\left\lfloor\left(\sqrt{9+8K}+1\right)/2\right\rfloor\right\}.$ (5)
@@
3.3 A Recursive Regression Perspective
Until now, our discussion has proceeded horizontally, framing the manifold dimension estimation problem as a model comparison task across different candidate dimensions $j$. A vertical examination of individual dimensions, equipped with specific model specifications, can provide complementary insights.
The local PCA procedure is typically regarded as a direct estimator of the tangent space. However, it also fits naturally within our regression framework. In particular, for $j=d$, if polynomial functions of increasing orders are adopted as models in $G_{d}$, then the three estimators—Local PCA, QE, and TLS—can all be interpreted as stages of a unified recursive regression scheme, where each stage fits a higher-order term of the local representation function based on the residuals from the previous step. We present this alternative view at the population level, which can be useful as it provides a potential pathway for statistical analysis of these methods.
Let $\bm{x}_{0}$ be a point on the manifold, and assume its neighborhood $\mathcal{N}_{\bm{x}_{0}}$ can be represented as the graph of a smooth function $g:\mathbb{R}^{d}\to\mathbb{R}^{p-d}$. Denote by $T:\mathbb{R}^{p}\to\mathbb{R}^{p}$ the linear change-of-coordinate mapping such that
$T(\bm{x}-\bm{x}_{0})=(\bm{x}_{1:d}^{\prime},\,g(\bm{x}_{1:d}^{\prime})),\qquad\forall\,\bm{x}\in\mathcal{N}_{\bm{x}_{0}}.$
This coordinate transform provides the starting point of the recursion. It aligns the local coordinates with the manifold geometry so that subsequent regressions can be expressed in these coordinates.
Stage 0: Estimating the coordinate transform
The first step is to estimate $T$ from data in the neighborhood of $\bm{x}_{0}$. Since $\bm{x}_{1:d}^{\prime}$ corresponds to the tangent space approximation around $\bm{x}_{0}$, at the population level, $T$ should be defined as the solution of
$\arg\min_{T\in\mathbb{R}^{p\times p}}\mathbb{E}\!\left[\|\bm{X}-P_{d}T(\bm{X}-\bm{x}_{0})\|^{2}\mid\bm{X}\in\mathcal{N}_{\bm{x}_{0}}\right],$
where $P_{d}:\mathbb{R}^{p}\to\mathbb{R}^{p}$ retains the first $d$ components of its input and makes the remaining 0.
When the conditional distribution of $\bm{X}$ around $\bm{x}_{0}$ is approximately uniform or at least isotropic, the solution should coincide with the population PCA, whose columns are the eigenvectors of the local covariance matrix
$\mathbb{E}_{\bm{x}_{0}}\!\left[(\bm{X}-\mathbb{E}_{\bm{x}_{0}}\bm{X})(\bm{X}-\mathbb{E}_{\bm{x}_{0}}\bm{X})^{\top}\right],$
where $\mathbb{E}_{\bm{x}_{0}}[\cdot]$ denotes conditional expectation given $\bm{X}\in\mathcal{N}_{\bm{x}_{0}}$. This provides a theoretical justification for our use of PCA coordinates to approximate the theoretical new coordinates. Empirically, this corresponds to setting $\bm{x}_{0}=\bm{x}_{k}$ and replacing population quantities with their sample analogues.
Stage 1: Linear model (no parameters to estimate)
After transforming the data via the estimated $\hat{T}$, we obtain new coordinates $\hat{\bm{X}}^{\prime}=\hat{T}(\bm{X}-\bm{x}_{0})$. At order $m=1$, the theoretical model is $p_{1}(\bm{x}_{1:d}^{\prime})=0$, which is fully known. Thus, no regression is needed at this stage (i.e., $\hat{p}_{1}=0$) —the linear structure has already been captured by the estimated tangent directions. If we stop here, we end up with Local PCA.
Stage 2: Quadratic model
The next step models the nonlinear geometry by fitting a quadratic map in the normal directions. Let
$p_{2\ell}(\bm{x}_{1:d}^{\prime})=\bm{x}_{1:d}^{\prime\top}Q_{\ell}\bm{x}_{1:d}^{\prime},\quad\ell=1,\ldots,p-d,$
where $Q_{\ell}=\nabla^{2}g_{\ell}(\bm{0})$. Using $\hat{\bm{X}}^{\prime}$ from the previous step, the goal is to estimate $Q_{1},\ldots,Q_{p-d}$ by regressing the remaining pdp-d coordinates on quadratic functions of the first $d$ coordinates. The population-level optimization problem is
$\arg\min_{Q_{\ell}\in\mathbb{R}^{d\times d}}\mathbb{E}\!\left[\sum_{\ell=1}^{p-d}\|\hat{X}_{\ell+d}^{\prime}-\hat{\bm{X}}_{1:d}^{\prime\top}Q_{\ell}\hat{\bm{X}}_{1:d}^{\prime}\|^{2}\mid\bm{X}\in\mathcal{N}_{\bm{x}_{0}}\right].$
Importantly, we can view this regression as being performed on the residuals from the previous stage,
$\hat{\bm{X}}_{(d+1):p}^{\prime}=\hat{\bm{X}}_{(d+1):p}^{\prime}-\hat{p}_{1}(\hat{\bm{X}}_{1:d}^{\prime}),$
which formalizes the recursive nature of the procedure.
Our QE and TLS estimators correspond to the special case of retaining only $\ell=1$, avoiding the combinatorial growth in the number of parameters. The former employs ordinary least squares, while the latter uses total least squares. Unlike the $m=1$ stage, the population solution here is not analytically tractable.
Stage 3 and beyond
In principle, one could continue recursively, fitting parameters of a third-order polynomial using the residuals from the quadratic stage,
$\hat{\bm{X}}_{(d+1):p}^{\prime}-\hat{p}_{1}(\hat{\bm{X}}_{1:d}^{\prime})-\hat{p}_{2}(\hat{\bm{X}}_{1:d}^{\prime}).$
However, the number of parameters increases rapidly even at $m=2$. Even accounting for symmetry, the number of coefficients corresponding to the third-order derivatives of $g$ is on the order of $p\times d^{3}/3!$, making accurate estimation statistically infeasible with local samples. In practice, we almost never go beyond a second-order model.